joe can paint a room in 6 hours less than billy. if they can paint the room in 4 hours working together, how long would it take each to paint the room working alone?
can you please solve this using the quadratic equations?? thanks!!

Answer by checkley71(8403)

X(X-6)/[X+(X-6)]=4 NOW CROSS MULTIPLY
X(X-6)=4[X+(X-6)]
X^2-6X=4X+4X-24
X^2-6X-8X+24=0
X^2-14X+24=0
(X-2)(X-12)=0
X-12=0
X=12 HOURS FOR BILLY TO PAINT THE ROOM ALONE.
12-6=6 HOURS FOR JOE TO PAINT THE ROOM.
PROOF:
(12*6)/(12+6)=4
72/18=4
4=4